OpenStudy (anonymous):

simplify cos(2*tan^-1(x))

OpenStudy (anonymous):

\[\cos(2\theta)=2\cos^2(\theta)-1\] so \[\cos(2\tan^{-1}(x))=2\cos^2(\tan^{-1}(x))-1\] is a good place to start

OpenStudy (anonymous):

all you need now is an expression for \[\cos(\tan^{-1}(x))\]

OpenStudy (anonymous):

thank you! I t think I know how to solve it now!

OpenStudy (anonymous):

yw